OpenStudy (anonymous):

Stats Problem Help please Using Composite Blood Samples When doing blood testing for HIV infections, the procedure can be made more efficient and less expensive by combining samples of blood specimens. If samples from three people are combined and the mixture tests negative, we know that all three individual samples are negative. Find the probability of a positive result for three samples combined into one mixture, assuming the proba- bility of an individual blood sample testing positive is 0.1 (the probability for the “at- risk” population, based on data from the New York State Health

OpenStudy (paxpolaris):

the mixture will test negative ONLY if all 3 samples are negative. the probability that all 3 samples are negative = 0.9*0.9*0.9

OpenStudy (perl):

the question asks, probability of testing positive

OpenStudy (paxpolaris):

=0.729 so probability that mixture will test positive = 1-0.729 = 0.271

OpenStudy (anonymous):

oh yes you are right! Pax,

OpenStudy (anonymous):

Thank you all!

OpenStudy (perl):

P (positive ) = P ( at least one of the 3 blood samples is positive) = 1 - P (all three are negative) by complement rule = 1 - (0.9)^3